Two APs have the same common difference .The difference between the 100th term is 100.what is the difference beteen the 1000th term?

Grade:10

1 Answers

Arun
6 years ago
Let two A.P's have the first terms a , a' and common difference d.
Given that difference between their 100th  term is 100.
then a + 99d - a' - 99d = 100
⇒ a - a' = 100.
Now difference between 1000 th term is
⇒ a + 999d - a' - 999d
= a - a'
= 100
∴ difference between 1000 th term is 100.